# Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

### v1

@@ 4.5 Iterative Analysis
Table 2: Ablation studies of TimeXL on real-world multi-modal time series datasets, evaluated by F1 score.
@@
Encoder Multi-modal 0.674 0.619 0.934 0.937
LLM Time(PromptCast) 0.499 0.418 0.727 0.696
Text 0.645 0.496 0.974 0.901
@@
4.6 Ablation Studies
In this subsection, we present the component ablations of TimeXL, as shown in Table 2, where we have several observations. Firstly, the performance of prediction LLM with text is better than PromptCast (Xue & Salim, 2023), which highlights the importance of contextual information for LLM in a zero-shot prediction scenario. Furthermore, the text prototypes consistently improve the predictive performance of LLM, underscoring the effectiveness of explainable artifacts from the multi-modal encoder, in terms of providing relevant contextual guidance. In addition, the fusion of prediction LLM and multi-modal encoder further boosts the predictive performance that surpasses the best of both multi-modal encoder and prediction LLM. These observations demonstrate the advantage of our framework synergizing the time series model and LLM for mutually augmented prediction. In Appendix B, full results (F1 and AUC) of TimeXL component ablation are provided in Table 4, and other ablations are provided in Figures 67.
@@ 6 Acknowledgments
7 Impact Statement
This work presents significant advancements in explainable multi-modal time series prediction by integrating time series encoders with large language model-based agents. The broader impact of this work is multifaceted. It has the potential to support high-stakes decision-making in domains such as finance and healthcare by delivering more accurate predictions accompanied by reliable case-based explanations that lead to more robust analyses. No ethical concerns must be considered in our work. The social impact is substantial as it provides a new paradigm for analyzing real-world multi-modal time series data through the integration of emerging AI tools like language agents.
@@ A.1 Dataset Statistics
In this subsection, we provide more details of the real-world datasets we used for the experiments. The data statistics are summarized in Table 3, including the meta information (e.g., domain resolution, duration of real-world time series records), the number of channels and timesteps and so on. We used the weather and healthcare datasets in TimeCAP (Lee et al., 2025), and the finance dataset in (Lee et al., 2024).
@@ A.2 Hyperparameters
First, we provide the hyperparameters of baseline methods. Unless otherwise specified, we used the default hyperparameters from the Time Series Library (Wu et al., 2023). For LLMTime (Gruver et al., 2023), OFA (Zhou et al., 2023), Time-LLM (Jin et al., 2024), TimeCMA (Liu et al., 2025), TimeCAP (Lee et al., 2025), we use their own implementations. For all methods, the dropout rate $\in\{0.0,0.1,0.2\}$, learning rate $\in\{0.0001,0.0003,0.001\}$. For transformer-based and LLM fine-tuning methods (Wu et al., 2021; Zhang & Yan, 2023; Liu et al., 2024c; Nie et al., 2023; Liu et al., 2025; Jin et al., 2024), the number of attention layers $\in\{1,2\}$, the number of attention heads $\in\{4,8,16\}$. For Dlinear (Zeng et al., 2023), moving average $\in\{3,5\}$. For TimesNet (Wu et al., 2023) the number of layers $\in\{1,2\}$. For PatchTST and MM-PatchTST, the patch size $\in\{3,5\}$ for the finance dataset.
Next, we provide the hyperparameters of TimeXL. The numbers of time series prototypes and text prototypes are $k\in\{5,10,15,20\}$ and $k^{\prime}\in[5,10]$, respectively. The hyperparameters controlling regularization strengths are $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}\in[0.1,0.3]$ with interval 0.05 for individual modality, $d_{\mathrm{min}}\in\{1.0,1.5,2.0\}$ for time series, $d_{\mathrm{min}}\in\{3.0,3.5,4.0\}$ for text. Learning rate for multi-modal encoder $\in\{0.0001,0.0003,0.001\}$ The number of case-based explanations fed to prediction LLM $\omega\in\{3,5,8,10\}$.
@@ A.3 Large Language Model
We employed the gpt-4o-2024-08-06 version for GPT-4o in OpenAI API. We use the parameters max_tokens=2048, top_p=1, and temperature=0.7 for content generation (self-reflection and text refinement), and 0.3 for prediction.
@@ Appendix B More Ablation Studies
Here we present the full results of TimeXL component ablations in Table 4. In addition to F1 scores, the results of AUC scores consistently demonstrate the importance of contextual information, the effectiveness of prototype from the multi-modal encoder, as well as the advantage of prediction fusion.
We also provide an ablation study on the learning objectives in the TimeXL encoder. The results clearly show that the full objective consistently achieves the best encoder prediction performance, highlighting the necessity of regularization terms that enhance the interpretability of multi-modal prototypes. The clustering ($\lambda_{1}$) and evidencing ($\lambda_{2}$) objectives also play a crucial role in accurate prediction: the clustering term ensures distinguishable prototypes across different classes, while the evidencing term ensures accurate projection onto training data.
Moreover, we assess how the number of matched case-based explanations enhances the prediction LLM, as shown in Figure 7. We conduct experiments on weather and finance datasets, demonstrating that incorporating more relevant case-based explanations consistently improves prediction performance. This further highlights the effectiveness of explainable artifacts in providing meaningful contextual guidance.
Figure 6: Ablation study of TimeXL encoder learning objectives.
Table 4: Ablation studies of TimeXL on real-world multi-modal time series datasets.
@@
Encoder Multi-modal 0.674 0.767 0.619 0.791 0.934 0.974 0.937 0.988
LLM Time(PromptCast) 0.499 0.365 0.418 0.607 0.727 0.768 0.696 0.871
Text 0.645 0.724 0.496 0.627 0.974 0.967 0.901 0.969
@@ C.1 Multi-modal Prototypes for All Datasets
We present the learned multi-modal prototypes across datasets including Weather (Figure 8), Healthcare (Test-Positive) (Figure 9), and Healthcare (Mortality) (Figure 10). It is noticeable that the prototypes from both modalities align well with real-world ground truth scenarios, ensuring faithful explanations and enhancing LLM predictions.
C.2 Case-based Reasoning Example on Finance
@@ Appendix D Designed Prompts for Experiments
In this section, we provide our prompts for prediction LLM in Figure 12 (and a text-only variant for comparison, in Figure 13), reflection LLM in Figures 1415 16, as well as refinement LLM in Figure 17. Note that we adopt a generate-update-summarize strategy to effectively capture the reflective thoughts from training samples with class imbalances, which is more structured and scalable. We make the whole training texts into batches. First, the reflection LLM generates the initial reflection (Figure 14) by extracting key insights from class-specific summaries, highlighting text patterns that contribute to correct and incorrect predictions. Next, it updates the reflection (Figure 15) by incorporating new training data, ensuring incremental and context-aware refinements. Finally, it summarizes multiple reflections from each class (in Figure 16) into a comprehensive guideline for downstream text refinement. This strategy consolidates knowledge from correct predictions while learning from incorrect ones, akin to the training process of deep models.
@@ Appendix F TimeXL for Regression-based Prediction: A Finance Demonstration
We conduct the experiments on the same finance dataset, where the target value is the raw material stock price. The performance of all baselines and TimeXL is presented in Table 5. Consistent with the classification setting, TimeXL achieves the best results, and the multi-modal variants of state-of-the-art baselines (MM-iTransformer and MM-PatchTSTS) benefit from incorporating additional text data. This observation is further supported by the ablation study in Table 6. We also note that the text modality offers complementary information for LLMs, enhancing both trend and price predictions. Furthermore, the identified prototypes provide contextual guidance, leading to additional performance gains. Overall, TimeXL outperforms all variants, underscoring the effectiveness of mutually augmented prediction.
@@
Figure 22: Prompt for regression tasks.
@@
DLinear (Zeng et al., 2023) 7.871 6.400 4.727
Autoformer (Wu et al., 2021) 7.215 5.680 4.263
Crossformer (Zhang & Yan, 2023) 7.205 5.313 3.808
TimesNet (Wu et al., 2023) 6.978 4.928 3.512
iTransformer (Liu et al., 2024c) 5.877 4.023 2.863
TSMixer (Chen et al., 2023a) 7.447 5.509 3.911
FreTS (Yi et al., 2024) 7.098 4.886 3.460
PatchTST (Nie et al., 2023) 5.676 4.042 2.853
LLMTime (Gruver et al., 2023) 11.545 5.300 3.774
PromptCast (Xue & Salim, 2023) 4.728 3.227 2.306
OFA (Zhou et al., 2023) 6.906 4.862 3.463
Time-LLM (Jin et al., 2024) 6.396 4.534 3.238
TimeCMA (Liu et al., 2025) 7.187 5.083 3.620
MM-iTransformer (Liu et al., 2024a) 5.454 3.789 2.687
MM-PatchTST (Liu et al., 2024a) 5.117 3.493 2.491
TimeXL 4.224 2.883 2.069
Table 6: Performance of TimeXL for regression tasks.
Variants F1 AUC RMSE MAE MAPE(%)
Multi-modal Encoder 0.610 0.792 4.281 2.926 2.099
Time(PromptCast) 0.418 0.607 4.728 3.227 2.306
Text + Time 0.520 0.651 4.848 3.167 2.238
Text + Time + Prototype 0.585 0.697 4.478 3.067 2.192
TimeXL 0.622 0.808 4.224 2.883 2.069

### v2

@@ 4.5 Iterative Analysis
4.6 Ablation Studies
Table 2: Ablation studies on real-world multi-modal time series datasets, evaluated by F1 score.
@@ 4.6 Ablation Studies
Encoder Time Series 0.602 0.585 0.889 0.861
Text 0.567 0.472 0.871 0.840
Multi-modal 0.674 0.619 0.934 0.937
LLM Text 0.645 0.496 0.974 0.901
@@
In this subsection, we present the component ablations of TimeXL, as shown in Table 2, where the texts are refined using the reflections with the best validation performance. Firstly, the multi-modal encoder consistently outperforms single-modality variants across all datasets, highlighting the benefit of integrating time series and textual modalities. Secondly, the prediction LLM exhibits better performance than text-only encoder, underscoring the advantage of LLM’s contextual understanding and reasoning. Furthermore, the text prototypes consistently improve the predictive performance of LLM, underscoring the effectiveness of explainable artifacts from the multi-modal encoder, in terms of providing relevant contextual guidance. Finally, the fusion of prediction LLM and multi-modal encoder further boosts the predictive performance that surpasses the best of both multi-modal encoder and prediction LLM. These observations demonstrate the advantage of our framework synergizing the time series model and LLM for mutually augmented prediction. In Appendix B, full results (F1 and AUC) of TimeXL component ablation are provided in Table 4, other ablations and studies (base LLM, quantitative evaluation of case-based explanations, learning objectives) are provided in Table 5 and Figures 678.
@@ A.1 Dataset Statistics
In this subsection, we provide more details of the real-world datasets we used for the experiments. The data statistics are summarized in Table 3, including the meta information (e.g., domain resolution, duration of real-world time series records), the number of channels and timesteps and so on. We used the weather and healthcare datasets in TimeCAP [81], and the finance dataset in [13].
@@ A.2 Hyperparameters
First, we provide the hyperparameters of baseline methods. Unless otherwise specified, we used the default hyperparameters from the Time Series Library (TSLib) [96]. For LLMTime [101], OFA [66], Time-LLM [69], TimeCMA [72], TimeCAP [81], FSCA [73], we use their own implementations. For all methods, the dropout rate $\in\{0.0,0.1,0.2\}$, learning rate $\in\{0.0001,0.0003,0.001\}$. For transformer-based and LLM fine-tuning methods [94, 95, 97, 5, 72, 69], the number of attention layers $\in\{1,2\}$, the number of attention heads $\in\{4,8,16\}$. For Dlinear [93], moving average $\in\{3,5\}$. For TimesNet [96] the number of layers $\in\{1,2\}$. For PatchTST [5], MM-PatchTST [47] and FSCA [73], the patch size $\in\{3,5\}$ for the finance dataset. For TimeCAP [81], we use the encoder with the prediction LLM, given the inherent multi-modal time series input. Next, we provide the hyperparameters of TimeXL. The numbers of time series prototypes and text prototypes are $k\in\{5,10,15,20\}$ and $k^{\prime}\in[5,10]$, respectively. The hyperparameters controlling regularization strengths are $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}\in[0.1,0.3]$ with interval 0.05 for individual modality, $d_{\mathrm{min}}\in\{1.0,1.5,2.0\}$ for time series, $d_{\mathrm{min}}\in\{3.0,3.5,4.0\}$ for text. Learning rate for multi-modal encoder $\in\{0.0001,0.0003,0.001\}$, using Adam [102] as the optimizer. The number of case-based explanations fed to prediction LLM $\omega\in\{3,5,8,10\}$. All results are reported as the average of five experimental runs.
@@ A.3 Large Language Model
We employed the gpt-4o-2024-08-06 version for GPT-4o in OpenAI API by default. We use the parameters max_tokens=2048, top_p=1, and temperature=0.7 for content generation (self-reflection and text refinement), and 0.3 for prediction. We keep the same setting for Gemini-2.0-Flash and GPT-4o-mini due to the best empirical performance.
@@ Appendix B More Ablation Studies
Table 4: Ablation studies of TimeXL on multi-modal time series datasets.
@@
Encoder Time Series 0.602 0.691 0.585 0.751 0.889 0.957 0.861 0.965
Text 0.567 0.658 0.472 0.636 0.871 0.941 0.840 0.887
Multi-modal 0.674 0.767 0.619 0.791 0.934 0.974 0.937 0.988
LLM Text 0.645 0.724 0.496 0.627 0.974 0.967 0.901 0.969
@@
Component ablations of TimeXL: Here we present the full results of TimeXL component ablations in Table 4. In addition to F1 scores, the results of AUC scores consistently demonstrate the importance of contextual information, the effectiveness of prototype from the multi-modal encoder, as well as the advantage of prediction fusion.
Ablations of learning objectives: We provide an ablation study on the learning objectives of TimeXL encoder, as shown in Figure 6. The results clearly show that the full objective consistently achieves the best encoder prediction performance, highlighting the necessity of regularization terms that enhance the interpretability of multi-modal prototypes. The clustering ($\lambda_{1}$) and evidencing ($\lambda_{2}$) objectives also play a crucial role in accurate prediction: the clustering term ensures distinguishable prototypes across different classes, while the evidencing term ensures accurate projection onto training data.
Effect of case-based explanations on LLM predictions and other quality assessments: Moreover, we assess how the number of allocated case-based explanations enhances the prediction LLM, as shown in Figure 7. We conduct experiments on weather and finance datasets, demonstrating that incorporating more relevant case-based explanations consistently improves prediction performance. This trend highlights the effectiveness of explainable artifacts in providing meaningful contextual guidance. Importantly, this analysis also serves as a proxy for quantitatively evaluating explanation quality, where retrieving higher-quality case-based examples tends to yield greater performance gains. Beyond this quantitative assessment, alternative strategies can be employed to evaluate explanation quality more explicitly, for example, using LLM as a judge [103] to rate the relevance and helpfulness of explanations in a human-understandable manner, with respect to ground truth.
Base LLM F1 AUC
GPT-4o-mini 0.932 0.981
Gemini-2.0-Flash 0.937 0.983
GPT-4o (default) 0.987 0.996
Table 5: Performance of different base LLMs on Healthcare(Test-Positive).
Effect of different base LLMs: To explore the effect of different base LLMs (Gemini-2.0 Flash and GPT-4o-mini), we provide the experiment results and analysis on Healthcare (Test-Positive) data. For fair comparisons, we used the same prompts and ran the same number of iterations and followed the same settings detailed in Appendix A.2,  A.3, and  A.4. The results are shown in Table 5. It can be observed that GPT-4o clearly outperforms both GPT-4o-mini and Gemini-2.0-Flash, due to its better reasoning and contextual understanding capabilities (larger model size & pre-training corpus). Both GPT-4o-mini and Gemini-2.0-Flash are still competitive compared with baselines listed in Table 1. It reveals the impact of the LLM capability on the effectiveness of our framework, especially in tasks demanding real-world context understanding. We also provide the plot of iterative analysis in Figure 8, where we can also observe the performance improvement over iterations for different base LLMs.
Figure 6: Ablation study of TimeXL encoder learning objectives.
Figure 8: Iterative study of different base LLMs on Healthcare (Test-Positive) dataset.
@@ C.1 Multi-modal Prototypes for All Datasets
We present the learned multi-modal prototypes across datasets including Weather (Figure 9), Finance (Figure 10), Healthcare (Test-Positive) (Figure 11), and Healthcare (Mortality) (Figure 12). It is noticeable that the prototypes from both modalities align well with real-world ground truth scenarios, ensuring faithful explanations and enhancing LLM predictions.
C.2 Case-based Reasoning Example on Finance Dataset
@@
Figure 10: Key multi-modal prototypes learned from Finance dataset. Each row in the figure represents a time series prototype.
@@ Appendix D Designed Prompts for Experiments
In this section, we provide our prompts for prediction LLM in Figure 14 (and a text-only variant for comparison, in Figure 15), reflection LLM in Figures 1617 18, as well as refinement LLM in Figure 19. Note that we adopt a generate-update-summarize strategy to effectively capture the reflective thoughts from training samples with class imbalances, which is more structured and scalable. We make the whole training texts into batches. First, the reflection LLM generates the initial reflection (Figure 16) by extracting key insights from class-specific summaries, highlighting text patterns that contribute to correct and incorrect predictions. Next, it updates the reflection (Figure 17) by incorporating new training data, ensuring incremental and context-aware refinements. Finally, it summarizes multiple reflections from each class (in Figure 18) into a comprehensive guideline for downstream text refinement. This strategy consolidates knowledge from correct predictions while learning from incorrect ones, akin to the training process of deep models.
@@ Appendix E Reflection Reports for Text Refinement
Figure 21: Reflection summary for text refinement of Finance dataset.
@@ Appendix F TimeXL for Regression-based Prediction: A Finance Demonstration
Figure 25: Prompt for prediction LLM in regression tasks.
@@
DLinear [93] 7.871 6.400 4.727
Autoformer [94] 7.215 5.680 4.263
Crossformer [95] 7.205 5.313 3.808
TimesNet [96] 6.978 4.928 3.512
iTransformer [97] 5.877 4.023 2.863
TSMixer [98] 7.447 5.509 3.911
TimeMixer [99] 6.651 4.703 3.349
FreTS [100] 7.098 4.886 3.460
PatchTST [5] 5.676 4.042 2.853
LLMTime [101] 11.545 5.300 3.774
PromptCast [79] 4.728 3.227 2.306
OFA [66] 6.906 4.862 3.463
Time-LLM [69] 6.396 4.534 3.238
TimeCMA [72] 7.187 5.083 3.620
FSCA [73] 5.511 3.873 2.720
MM-iTransformer [47] 5.454 3.789 2.687
MM-PatchTST [47] 5.117 3.493 2.491
TimeCAP [81] 4.456 3.088 2.196
TimeXL 4.161 2.844 2.035
We conduct the experiments on the same finance dataset, where the target value is the raw material stock price. The performance of all baselines and TimeXL is presented in Table 6. Consistent with the classification setting, TimeXL achieves the best results, and the multi-modal variants of state-of-the-art baselines (MM-iTransformer and MM-PatchTST) benefit from incorporating additional text data. This observation is further supported by the ablation study in Table 7. We also note that the text modality offers complementary information for LLMs, enhancing numerical price predictions. Furthermore, the identified prototypes provide contextual guidance, leading to additional performance gains. Overall, TimeXL outperforms all variants, underscoring the effectiveness of mutually augmented prediction. We also provide an iteration analysis to show the effectiveness of reflective and refinement process, as shown in Table 8. The prediction performance quickly improves and stabilizes over iterations, which underscores the alternation steps between predictions and reflective refinements. Additionally, we provide a visualization in Figure 26, which compares the predicted trend and stock price values, with ground truths for reference.
Table 7: Ablation studies of TimeXL for regression tasks.
Variants RMSE MAE MAPE(%)
Time Series Prototype-based Encoder 4.287 3.001 2.140
Multi-modal Prototype-based Encoder 4.198 2.891 2.064
Prediction LLM - Time Series + Text 4.600 3.121 2.226
Prediction LLM - Time Series + Text + Prototype 4.352 3.003 2.165
TimeXL 4.161 2.844 2.035
Table 8: Iterative analysis of TimeXL for regression tasks.
Iteration RMSE MAE MAPE(%)
Original 4.344 2.951 2.103
1 4.224 2.883 2.069
2 4.161 2.844 2.035
3 4.174 2.849 2.036
Figure 26: Visualization of TimeXL’s numerical and trend predictions on Finance dataset.
Finally, we provide a case study on the testing set of finance data that demonstrate the effectiveness of TimeXL’s reasoning process for regression tasks. We present the top three most similar prototypes allocated for both time series and text modalities, where we have two key observations. 1) Even if the trend predictions align correctly in both original and refined texts, the prototype allocation varies regarding market sentiments. In the original texts, prototypes from the neutral class are also allocated, highlighting a mixed sentiment and potential fluctuations, consequently deteriorating numerical prediction accuracy from both the encoder and the LLM. In the refined texts, prototypes instead capture the underlying bearish market sentiment for iron ore, emphasizing decreased demand and weakened market conditions, effectively aligning bearish market sentiment with numerical predictions. 2) The allocated top similar time series prototypes also exhibit highly similar trends compared to the original input series, offering robust reference points for numerical prediction. To quantitatively assess it, we aggregate the retrieved prototypes using their similarity scores and compute the correlation between the input time series and the aggregated prototype series. The observed strong correlations underscores TimeXL’s efficacy in capturing intrinsic temporal patterns and providing predictive insights into future numerical stock values based on similar historical patterns.
Figure 27: Multi-modal case-based reasoning example on Finance dataset for regression tasks.
Appendix G More Detailed Discussions of Related Work
Due to space limitations, we provide additional details on related methods here, expanding on the overview in the main text.
G.1 Multi-modal Time Series Analysis.
In recent years, multi-modal time series analysis has gained significant traction in diverse domains such as finance, healthcare, environmental sciences, and industry [11, 38, 12, 39, 16, 14]. Multiple approaches have been proposed to model interactions across different modalities, such as multi-modal fusion and cross-modal alignment and so on. Multi-modal fusion captures complementary information to enhance time series modeling, which are commonly implemented via addition [47] and concatenation [40, 70, 47, 41] at the multiple levels. For example, Liu et al. [47] fuse predictions from both time series and text modalities via learnable linear weighted addition, further enhancing the predictive performance of state-of-the-art forecasters. In parallel, cross-modal alignment aims to capture the association between time series and other modalities for downstream tasks, where common techniques include attention mechanism [13, 42, 104, 43], gating functions [44], and contrastive learning [45, 46]. For instance, Bamford et al. [17] align time series and text in a shared latent space using deep encoders, enabling retrieval of specific time series sequences based on textual queries. In addition, Zheng et al. [45] perform causal structure learning in multi-modal time series by separating modality-invariant and modality-specific components through contrastive learning, enabling root cause analysis. Besides method developments, recent research efforts have also advanced the field through comprehensive benchmarks and studies [47, 48, 49, 50, 51], highlighting the incorporation of new modalities to boost task performance. Meanwhile, others explore representing time series as other modalities (e.g., image, text, table, graph) [52, 53, 54, 55, 105] or generating synthetic modalities [44, 56] to facilitate downstream tasks. Nevertheless, these studies tend to focus primarily on improving numerical accuracy. The deeper and tractable rationale behind why or how the textual or other contextual signals influence time series outcomes remains underexplored.
G.2 Time Series Explanation
Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based explanations leverage saliency methods to highlight important features at different time steps [57, 58], where some methods explicitly incorporate temporal structures into models and objectives [59, 60]. Meanwhile, surrogate approaches also offer global or local explanations by evaluating the importance of time series features with respect to prediction. These methods include the the generalization of Shapley value to time series [61], the formulation of model behavior consistency via self-supervised learning [62], and the usage of information-theoretic guided objectives for coherent explanations [63]. In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm [34, 35, 36, 37], which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
G.3 LLMs for Time Series Analysis
The rapid development of LLMs [18, 19, 20] has begun to inspire new directions in time series research [64, 65]. Many existing methods fine-tune pre-trained LLMs on time series tasks and achieving state-of-the-art results in forecasting, classification, and beyond [66, 67, 68]. The textual data (e.g., domain instructions, metadata, and dataset summaries) are often encoded as prefix embeddings to enrich time series representations [69, 70, 71, 72, 73, 74]. These techniques also contribute to the emergence of time series foundation models [68, 75, 76, 77, 78]. An alternative line of research leverages the zero-shot or few-shot capabilities of LLMs by directly prompting pre-trained language models with text-converted time series [79] or context-laden prompts representing domain knowledge [30, 32, 33], often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines [80, 31, 81, 82]. For instance, Wang et al. [82] construct a large news database for method development and perform text-to-text prediction by fine-tuning an LLM with dynamically selected news, which uses reflective selection logic to incorporate matched social events and thus augment numerical prediction. Building on this trend, recent works have explored time series reasoning with LLMs by framing tasks as natural-language problems [84], including time series understanding & question answering (e.g., trend and seasonal pattern recognition, similarity comparison, causality analysis, etc.[78, 86, 87, 83, 85, 88, 89], as well as language-guided inference for standard zero-shot time series tasks [83, 90, 85]. Compared to existing LLM-based methods, our approach provides a unique synergy between an explainable model and LLMs through iterative, grounded, and reflective interaction. It explicitly provides case-based explanations from both modalities, highlighting key time series patterns and refined textual segments that contribute to the prediction.
Appendix H Limitations
Our main focus has been on classification-based time series prediction to emphasize the explanation capabilities in real-world multi-modal applications, while we also provide a detailed regression-based prediction setting in the appendix. Extending our framework to broader regression-based tasks under the multi-modal contexts, such as long-term time series forecasting, and time series forecasting with domain shifts, remains an important direction for us to explore in the future work.
Appendix I Broader Impacts
Our paper presents advancements in explainable multi-modal time series prediction by integrating time series encoders with large language model-based agents. The broader impact of this work is multifaceted. It has the potential to support high-stakes decision-making in domains such as finance and healthcare by delivering more accurate predictions accompanied by reliable case-based explanations that lead to more robust analyses. The social impact of this work stems from its potential to provide a new paradigm for analyzing real-world multi-modal time series through the integration of emerging AI tools such as language agents. We emphasize the importance of responsible and ethical deployment of LLMs in time series analysis to ensure that such advancements lead to positive societal outcomes.